### v1

# SpotVLM: Cloud-edge Collaborative Real-time VLM based on Context Transfer

SpotVLM: Cloud-edge Collaborative Real-time VLM
based on Context Transfer
Chen Qian1, Xinran Yu1, Zewen Huang1, Danyang Li1, Qiang Ma1,
Fan Dang2, Xuan Ding1, Guangyong Shang3, Zheng Yang1
1 Abstract
Vision–Language Models (VLMs) are increasingly deployed in real-time applications such as autonomous driving and human–computer interaction, which demand fast and reliable responses based on accurate perception. To meet these requirements, existing systems commonly employ cloud–edge collaborative architectures, such as partitioned Large Vision-Language Models (LVLMs) or task offloading strategies between Large and Small Vision-Language Models (SVLMs). However, these methods fail to accommodate cloud latency fluctuations and overlook the full potential of delayed but accurate LVLM responses. In this work, we propose a novel cloud–edge collaborative paradigm for VLMs, termed Context Transfer, which treats the delayed outputs of LVLMs as historical context to provide real-time guidance for SVLMs inference. Based on this paradigm, we design SpotVLM, which incorporates both context replacement and visual focus modules to refine historical textual input and enhance visual grounding consistency. Extensive experiments on three real-time vision tasks across four datasets demonstrate the effectiveness of the proposed framework. The new paradigm lays the groundwork for more effective and latency-aware collaboration strategies in future VLM systems.
2 Introduction
Figure 1: Overview of representative deployment and inference strategies for VLMs.
VLMs have witnessed rapid advancements in recent years. These models typically consist of a visual encoder, a language decoder, and a cross-modal fusion module, enabling them to perform a wide range of tasks such as image captioning (Xue et al. 2025), visual question answering (Fang et al. 2025), and robotic perception (Zhou et al. 2025). Notably, LVLMs (Liu et al. 2025), empowered by large-scale model capacity and extensive training data, demonstrate superior generalization ability and reasoning performance across a broad range of downstream tasks. However, due to their substantial computational demands, these models are often deployed on the cloud servers, which inevitably involve network transmission, model inference, and request queuing delays, posing challenges for scenarios that require real-time responses within seconds.
To overcome these limitations, recent studies have proposed lightweight SVLMs for edge deployment (Yao et al. 2024; Rang et al. 2025), enabling low-latency, on-device inference suitable for real-time scenarios like autonomous driving (Lu et al. 2025; Sima et al. 2024; Tian et al. 2024), augmented reality (Pan and Liu 2025), and smart wearable devices (Huang et al. 2024). However, due to their limited model capacity, SVLMs still exhibit significantly lower inference accuracy than LVLMs (Zhao et al. 2025).
To enable efficient and accurate real-time visual reasoning, recent studies (Hu et al. 2024b; Li et al. 2025b; Hu et al. 2024a) have explored cloud–edge collaborative architectures. As illustrated in Fig.1, existing approaches fall into two categories: (i) Model partitioning, which splits the LVLM across the cloud and edge based on hardware constraints. For example, Distributed VLM (Li et al. 2025b) deploys the visual encoder on the edge and transmits extracted features to the cloud, where they are fused with text and processed by an LLM. (ii) Dynamic task offloading, which routes inputs based on task complexity and system conditions. For instance, ADAS (Hu et al. 2024b) offloads tasks to the cloud for LVLM inference by considering factors such as latency and quality of service (QoS).
Although such methods reduce cloud-side inference workload, they still face performance bottlenecks from service latency. In video streaming scenarios, accumulated delays often cause the outputs of LVLMs arriving several frames behind the current input (e.g., producing results for frame t–d while the system is processing frame t). Consequently, although these outputs are of high quality, they are temporally outdated and cannot be effectively utilized for real-time reasoning.
To further explore more effective collaboration paradigms between large and small VLMs, we analyze the underlying factors that contribute to their performance gap and derive the following insights:
$\bullet$ Utilization of High-Quality Textual Context. Historical context plays a crucial role in generating high-quality outputs. For example, Qwen-VL (Bai et al. 2023) adopts an accumulative input structure in multi-turn interactions: each round of input consists of the input and response from the previous round, along with the current query. This context is progressively extended until the window reaches the maximum memory or model-defined limit. Such accumulation allows the model to integrate information across turns, enhancing semantic consistency and reasoning depth. More importantly, the generated outputs are recursively fed into subsequent inputs, forming a performance-enhancing feedback loop. Although small models adopt similar input strategies, their limited generative capacity often results in lower-quality historical content, which negatively affects reasoning performance. Enhancing the quality of contextual history thus emerges as a key factor in improving the reasoning ability of small models.
$\bullet$ Grounding of Semantically Salient Image Regions. The ability to align visual content with textual descriptions is critical in vision-language reasoning. However, SVLMs often struggle in this aspect, particularly when dealing with fine-grained semantics or complex scenes. Prior studies (Zheng et al. 2025; Xu et al. 2025; Zhang et al. 2025a) have demonstrated that LVLMs, leveraging their grounding capabilities, can automatically crop semantically rich image regions as refined inputs—thereby enhancing the focus and semantic interpretability in visual reasoning. Guiding small models to focus on task-relevant image regions enhances their ability to interpret complex visual and linguistic cues, ultimately improving their reasoning performance.
Our Work: Given the inherent limitations of cloud-based LVLMs in terms of latency and response stability, we propose a key design principle in this work: Context Transfer. Instead of relying on the delayed outputs from LVLMs as real-time feedback, they are reused as contextual priors to enhance future reasoning in SVLMs.
Building on this principle, we introduce SpotVLM, which incorporates two collaborative modules that guide the SVLM along both the linguistic and visual dimensions to achieve more accurate multimodal reasoning.
Motivated by the strong impact of historical context quality on VLM inference, we design a Context Replacement Module (CRM). Specifically, It extracts key textual segments from high-quality LVLM responses, reorganizes them based on task context, and embeds them as contextual history to replace the original low-quality history. These informative exemplars implicitly guide the small model to focus on semantically relevant visual cues that were previously overlooked, thereby enhancing cross-modal association and reasoning performance.
To exploit the grounding capabilities of large models, we introduce the Visual Focus Module (VFM), which identifies semantically salient regions from historical images based on LVLM predictions. These regions are then used to guide the SVLM: non-essential patches are masked to reduce visual tokens and improve efficiency, while attention is focused on informative areas to enhance grounding and alignment. To adapt visual focus dynamically, the edge device performs feature matching between historical and current images. Semantically aligned regions are assigned higher weights, allowing the SVLM to inherit and shift attention toward task-relevant areas in the current frame.
By providing high-quality textual history and guiding attention to semantically relevant visual regions, these modules effectively leverage the delayed outputs of LVLMs to compensate for the limited reasoning capacity of SVLMs.
The contributions of this work are summarized as follows:
  • We propose a novel architectural design paradigm, termed Context Transfer, for real-time vision-language reasoning. Instead of relying on the final predictions from LVLMs, this paradigm treats their delayed outputs as contextual priors to guide SVLMs, enabling a new form of efficient large–small model collaboration.
  • Based on this principle, we design SpotVLM with two zero-shot collaborative modules: the Context Replacement Module and the Visual Focus Module, which guide the SVLMs to reuse high-quality outputs from the LVLMs along the linguistic and visual dimensions, respectively.
  • Extensive experiments on four public benchmarks validate the effectiveness of SpotVLM, showing consistent improvements over prior cloud–edge collaborative approaches.
3 Related Work
3.1 Cloud-Edge Collaborative VLMs
Figure 2: Overview of the proposed collaboration framework, SpotVLM. The system takes timestamped video streams as input and performs two parallel operations: uploading selected keyframes to the cloud-based LVLM for processing, and conducting local inference using the SVLM. Given potential cloud latency, delayed LVLM outputs are reused as historical context. These outputs guide the SVLM through two modules, Context Replacement and Visual Focus, to improve the quality of real-time predictions.
Driven by the growing demand for real-time VLM-based processing, applications such as autonomous driving (Lu et al. 2025) are shifting from modular to end-to-end pipelines (Shao et al. 2024; Li et al. 2025a), where latency sensitivity is critical and excessive delays can lead to system failure.
To enable fast and robust responses, cloud-edge collaborative architectures for VLMs have attracted increasing attention. Existing collaboration paradigms can be categorized into two types. The first category is model partitioning (Li et al. 2025b; Zhang et al. 2025b). To improve processing efficiency for LVLMs in the cloud, Distributed VLM (Li et al. 2025b) offload the visual encoder to the edge device, which encodes input images rapidly and transmits the feature embeddings to the cloud for processing and decoding. VaVLM (Zhang et al. 2025b) further reduces transmission overhead by uploading only the region-of-interest (ROI) of images. This parallel architecture entirely rely on stable and responsive cloud services, making it difficult to guarantee robust frame-wise inference in real-time settings. The second category involves dynamic task allocation to different models according to predefined offloading strategies. Some systems (Hu et al. 2024b) formulate the offloading decision as an optimization problem, incorporating multiple system-level factors such as Quality of Service (QoS) and latency. In addition, frameworks such as LAECIPS (Hu et al. 2024a) pretrain a dedicated task classifier to estimate task complexity and offload difficult samples to cloud-based large models. While these approaches leverage the complementary strengths of large and small models to reduce cloud-side computation, they overlook the continued guidance that high-quality LVLM responses can provide for subsequent inference. Recognizing this underexplored potential, we propose a real-time collaborative framework that integrates delayed LVLM outputs into SVLM context to enable continual refinement and robust inference.
3.2 Historical Context in VLMs
Context refers to the sequence of prior interactions retained to guide subsequent inferences, essential for maintaining coherence in multi-turn tasks. In VLMs, this mechanism is adapted from Large Language Models (LLMs), which use fixed-length context windows to preserve relevant token sequences and enhance reasoning consistency. Prior research on LLMs has shown that adjusting context improves model performance, as reflected in translation accuracy (Sung et al. 2024) and dialogue consistency (Wei et al. 2024). Similarly, VLMs leverage effective context preservation to significantly improve robotic navigation in embodied AI scenarios (Habibpour and Afghah 2025), reducing the risk of decision paralysis and oscillatory behavior. Building on this advantage, the proposed SpotVLM retain the output from LVLMs as high-quality context to guide the SVLMs, helping it focus on salient content and improve inference quality.
4 Architecture
Real-time vision-language reasoning tasks pose dual challenges in terms of processing speed and inference accuracy. To achieve robust and high-precision outputs, this work leverages the strengths of LVLMs in textual understanding and visual perception to assist SVLMs in real time. This section illustrates the proposed collaborative architecture using a multi-object recognition task in autonomous driving at a frame rate of 1 FPS as a representative example.
4.1 Workflow
The front-facing camera continuously captures frames and sends them to the edge device. For each timestamped frame, the edge decides whether to upload it to the cloud based on the latency of the previous cloud interaction. Each frame is also processed locally by the SVLM for real-time inference. For uploaded frames, LVLM responses are handled according to latency: results within 1 second are adopted and stored, while delayed outputs are retained for later use, and real-time results from the SVLM are adopted for the current frame.
As illustrated in Fig.2, the output returned by the LVLM contains two key components: (1) the response content to the given query, denoted as <answer>; and (2) the attended image regions during inference, indicated as <ROI>. These two outputs provide guidance for subsequent frames from textual and visual perspectives, respectively. In addition, the returned result includes the timestamp of the original frame, which allows the system to assess current cloud latency and determine the temporal relevance of the result for semantic guidance.
4.2 Context Replacement
In multi-turn vision-language reasoning tasks, VLMs typically take an image and a textual question from the user as the current input. To support continuous multimodal understanding, VLMs often retain part of the historical context during inference. This insight informs the design of two novel mechanisms for enabling textual-level collaboration across LVLMs and SVLMs.
As detailed in Sec.4.1, high‑quality outputs from the large model are continuously updated and stored in an edge‑side cache as $\mathcal{H}_{T-d}^{LVLM}$:
$\mathcal{H}_{T-d}^{LVLM}\leftarrow LVLM(I_{T-d},Q_{T-d}),$ (1)
where $I_{T-d}$ and $Q_{T-d}$ denote the image and question for Frame T-d, respectively. When processing frame T, the system compares the current timestamp with the most recent timestamp T-d associated with a returned LVLM result. If the delay is within a predefined threshold $\delta$, the SVLM replaces its original output for frame T-d with the response from LVLM, and incorporates it into the historical context to guide inference on frame T.
$if\quad\Delta t=T-(T-d)\leq\delta,\quad\mathcal{H}_{T-d}^{SVLM}\leftarrow\mathcal{H}_{T-d}^{LVLM}$ (2)
Given the temporal continuity and strong correlation between adjacent frames, providing more accurate historical information increases the likelihood of correctly activating relevant tokens and suppressing irrelevant ones, thereby improving the overall precision of outputs generated by the SVLM.
Moreover, instead of directly concatenating raw questions and answers as historical context, we design an in-context learning prompt template to compress and summarize dialogue history for the SVLM, and the result is denoted as follows:
$\tilde{\mathcal{H}}=Compress(\mathcal{H}_{\leq T-1}).$ (3)
Given the limited memory capacity of edge devices and the context window constraints of lightweight models, naively stacking multi-turn inputs may lead to token truncation. By compressing the historical content, our method not only reduces the number of irrelevant tokens but also enhances attention to the semantics of prior answers for the model.
To summarize, with the textual collaboration strategy in place, the SVLM receives an input structure at frame T that corresponds to the right side of the Context Replacement module in the figure.
$Input_{T}^{SVLM}=\{I_{T},Q_{T},\tilde{\mathcal{H}}\}$ (4)
This input incorporates distilled historical answers from the LVLM, enabling the SVLM to perform accurate and efficient semantic reasoning.
Figure 3: ROI-based region tracking via cosine similarity. Given ROIs of traffic cone and pedestrian from frame T–d, cosine similarity highlights related regions in frame T to guide the attention of the small model.
4.3 Visual Focus
LVLMs, equipped with strong grounding capabilities, identifies task-relevant image regions and transmits them to the edge-side SVLMs. Based on this, SVLMs can proactively focus on critical visual regions from two perspectives.
Grounding-Guided Visual Representation.
The ROIs returned by the LVLM guide the grounding of the SVLM on historical frames. Most VLMs, including the Qwen-2.5-VL-3B (Bai et al. 2025), use a Vision Transformer (ViT) that divides a resized input image into patches and encodes them into visual tokens (e.g., 1824 tokens for a 1600 × 900 image). However, many of these tokens are irrelevant to the reasoning task. To accelerate inference and enhance focus, ROIs are mapped to corresponding patches, and only task-relevant ones are selected. This selection is formally defined as:
$\mathcal{P}_{ROI}=\left\{p_{i}\in\mathcal{P}\mid p_{i}\cap ROI\neq\emptyset\right\},$ (5)
where $\mathcal{P}$ denotes the set of all image patches, and $p_{i}$ is a patch whose spatial overlap with the ROI is non-empty. In the case of multi-object recognition, the largest ROI is selected for each target category. For the case illustrated in the figure, the number of features selected for projection is reduced to 49, thereby reducing the computational cost.
nuScenes BDD100K MOT IPN Hand ACE
Micro-F1 Macro-F1 0-1 EM Micro-F1 Macro-F1 0-1 EM Acc. Acc.
SVLM Only (Local) 58.19 37.70 16.10 62.12 31.06 6.37 36.41 23.43
LVLM Only (Cloud) 45.76 32.04 20.75 49.65 25.85 20.28 23.91 61.14
Distributed VLM (Li et al. 2025b) 45.31 31.78 20.39 49.31 25.71 20.09 23.64 60.86
LAECIPS (Hu et al. 2024a) 59.84 39.64 24.29 62.47 32.14 21.05 39.40 48.71
ADAS (Hu et al. 2024b) 60.00 41.35 18.11 63.87 33.12 10.63 39.27 55.43
SpotVLM 69.96 48.19 27.32 71.19 36.26 23.88 45.11 62.75
Table 1: Performance comparison of different architectures across four datasets (values in %). All methods adopt Qwen-VL-2.5-3B as the SVLM and Qwen-VL-2.5-72B as the LVLM.
Grounding-Tracked Visual Representation.
The ROIs provided by the LVLM support cross-frame tracking, guiding the SVLM to focus on relevant regions in subsequent frames. Due to limited grounding capability, lightweight SVLMs may struggle to localize objects even when their categories are known. However, given the temporal continuity of video, ROI features from frame T-d can be matched to visual features in frame T. Let $\mathcal{P}_{T-d}^{(c)}$ denotes the set of patch indices for category $c$ on frame T-d, and $\mathbf{v}_{i}^{(T-d)}$ is the visual feature of the $i$-th patch. Thus, feature vector of object category $c$ on frame T-d is:
$\mathbf{f}^{(c)}=\frac{1}{|\mathcal{P}^{(c)}_{T-d}|}\sum_{i\in\mathcal{P}^{(c)}_{T-d}}\mathbf{v}^{(T-d)}_{i}$. (6)
We compute the cosine similarity between ROI features from frame T-d and the visual features of frame T, thereby obtaining similarity maps $S^{(c)}$ for each potential category $c$ in frame T.
$\mathbf{S}^{(c)}=\left[\frac{\mathbf{f}^{(c)}\cdot\mathbf{v}^{(T)}_{j}}{\|\mathbf{f}^{(c)}\|\cdot\|\mathbf{v}^{(T)}_{j}\|}\right]_{j=1}^{N},$ (7)
where $v^{(T)}_{j}$ denotes the feature of the $j$-th patch on frame T, and $N$ is the total number of patches. As shown in Fig.3, the LVLM accurately detects and localizes small objects such as traffic cones and pedestrians in frame T-d, and their mapped similarity heatmaps on frame T effectively indicate the potential locations after viewpoint shift, providing precise visual guidance for the SVLM. This cross-frame computation yields patch-level similarity scores can be aggregated into a unified similarity vector $s$, which can be defined as:
$s=\frac{1}{C}\sum_{c=1}^{C}S^{(c)}.$ (8)
Then $s$ can be used to construct the grounding-tracked visual weight $W_{s}$, formulated as follows:
$W_{s}=\mu\cdot\frac{1}{1+e^{-s}}+b,$ (9)
where $\mu$ and $b$ are tunable scaling parameters used to control the visual weights.
To further adapt the visual representation based on semantic associations, we introduce a gated token update mechanism that enables dynamic modulation of visual tokens. The update is formulated as:
$t\leftarrow(1-\alpha)\cdot t+\alpha\cdot W_{s}\cdot t.$ (10)
Here, $t$ denotes the original visual token, and $\alpha\in[0,1]$ is a control coefficient that balances the weights between known object categories and newly emerging targets. The dynamically predicted weight $W_{s}$ modulates the visual tokens based on feature similarity, enabling adaptive token refinement under the guidance of semantic consistency.
5 Experiments
5.1 Experiments Setting
Datasets.
We evaluate the proposed architecture on three real-time tasks.
Real-time multi-object recognition. (i) nuScenes (Caesar et al. 2019): We sample camera images at 1 FPS from this autonomous driving dataset. The processed data includes 150 scenes (20 seconds each) with 10 object categories. (ii) BDD100K (MOT) (Yu et al. 2020): We sample the driving video frame at 1 FPS. The processed data includes 200 scenes (40 seconds each) with 10 object categories.
Real-time gesture recognition. IPN Hands (Benitez-Garcia et al. 2021): A gesture recognition dataset with 14 gesture classes, sampled at 2 FPS from 52 close-range videos (avg. 120s each) to capture dynamic motion.
Real-time video frame caption. Actions for Cooking Eggs (ACE) (Shimada et al. 2012): A kitchen scene dataset with 8 cooking actions, sampled at 1 FPS from 10 videos (avg. 194s each).
Metrics.
We adopt commonly used evaluation metrics (Durand, Mehrasa, and Mori 2019; Nam et al. 2017; Huang et al. 2017) for multi-object recognition task, which are Micro-F1, Macro-F1 and 0-1 Exact Match (0-1 EM). Specifically, Micro-F1 reflects overall performance across all instances, and Macro-F1 captures class-wise performance equally. Additionally, 0-1 EM metric equals 1 only if the predicted result exactly matches the ground truth label for the current frame.
Following 0-1 EM, we model both gesture recognition and video frame captioning as frame-level single-label classification tasks, and use accuracy (Acc) as the evaluation metric.
Note that the evaluated tasks operate under real-time constraints (1–2 FPS). If the system fails to return a valid result within one second, the output for that frame is considered empty.
Figure 4: Qualitative Results. Illustration of input questions, ground-truth answers, and system predictions from four datasets.
Implementation Details.
In the Visual Focus Module, the parameters $\mu$, $b$ and $\alpha$ are set to 0.5, 0.75, and 0.1, respectively. To balance processing speed and model performance, we retain two previous frames as context. SpotVLM consistently treats the cloud feedback as the farthest frame in the historical context of SVLM.
In addition, to evaluate the effectiveness of different collaborative architectures, we use the official pretrained versions of both the LVLMs and SVLMs without any additional fine-tuning. During inference, the outputs and response latencies of the large model are recorded for each frame to ensure fair comparison of collaboration strategies. All experiments on the edge side are conducted using a GeForce RTX 4090 GPU with 24 GB of memory.
CRM GGVR GTVR Micro-F1 Macro-F1 0-1 EM
66.87 43.07 24.39
71.46 48.64 29.98
72.15 50.85 30.93
73.02 51.56 33.17
Table 2: Ablation Study. To assess the contribution of each component, we use full-frame upload and LVLM–SVLM parallel inference as the baseline, and evaluate performance improvements as modules are added incrementally.
SVLM LVLM Micro-F1 Macro-F1 0-1 EM
Baseline SpotVLM Baseline SpotVLM Baseline SpotVLM
Qwen2.5-VL-3B Qwen2.5-VL-72B 66.87 73.02 43.07 51.56 24.39 33.17
MiniCPM-o-2.6-int4 Qwen2.5-VL-72B 67.68 73.56 43.47 49.63 31.38 35.71
Qwen2.5-VL-3B GPT-4o 66.34 74.25 47.85 53.74 22.53 30.79
Table 3: Adaptability Study. Testing the SpotVLM architecture with various SVLM–LVLM combinations on nuScenes.
5.2 Overall Performance
To evaluate the effectiveness of different architectures, we assess six methods under a Wi-Fi network environment: locally deployed Qwen2.5‑VL‑3B (Bai et al. 2025), cloud-based Qwen2.5‑VL‑72B (Bai et al. 2025), Distributed VLM, LAECIPS, ADAS, and SpotVLM. The results are presented in Tab.1.
Due to cloud latency, methods that rely entirely on large models (e.g., Qwen2.5-VL-72B, Distributed VLM) yield lower F1 scores on real-time tasks than collaborative approaches with small models. However, the 0–1 EM of Qwen2.5-VL-72B remains higher than that of the small model alone, showing its superior accuracy when responses are timely. Additionally, Qwen2.5-VL-3B produces visual features larger than the original image, making Distributed VLM, which transmits features to the cloud, slightly less effective than running full inference on the cloud model.
Collaborative architectures like LAECIPS and SpotVLM perform consistently well across all four datasets. When cloud responses are timely, they effectively leverage large-model outputs, achieving better overall metrics than Qwen2.5-VL-3B.
ADAS adopts a latency-adaptive offloading strategy. In the first three datasets, long outputs from multi-object recognition and dual-image inputs in gesture recognition result in high cloud latency. Under these conditions, ADAS relies more on the small model, yielding higher F1 scores but lower 0–1 EM compared to LAECIPS. In the fourth dataset, shorter outputs and fewer images per inference lead to more timely cloud responses. ADAS thus relies more on the output of the LVLM. Meanwhile, LAECIPS depends more on the SVLM for simple samples in this dataset, resulting in slightly lower performance than ADAS.
To ensure a fair comparison across methods, we recorded the number of cloud service invocations for ADAS and LAECIPS, and used their average values to set the interval for SpotVLM. Benefiting from its ability to effectively leverage the high-quality outputs of the LVLM under latency constraints, SpotVLM consistently achieved the best results across all tasks and collaborative strategies.
To provide a more intuitive comparison, we select one representative example from each of the four datasets for qualitative analysis. As shown in the Fig.4, for the real-time multi-object recognition task, all three methods rely on the outputs of the SVLM, with SpotVLM achieving the best performance due to guidance from the historical results of the large model. In the gesture recognition task, as two images need to be transmitted and processed simultaneously, the cloud-based large model fails to return valid results in real time, leading to No response for ADAS. In contrast, SpotVLM leverages the history-enhanced small model to produce accurate classification results. For the video-frame captioning task, both ADAS and SpotVLM adopt the results quickly returned by the large model.
5.3 Ablation Study
We conduct a progressive ablation study on the nuScenes dataset to evaluate the individual and combined contributions of Contextual Replacement Module (CRM), Grounding-Guided Visual Representation (GGVR), and Grounding-Tracked Visual Representation (GTVR) in the SpotVLM framework. To isolate component effects, all image frames are uploaded, eliminating the influence of frame selection and upload strategies.
As shown in the Tab.2, the first row corresponds to the baseline where the large and small models operate in parallel—using the LVLM output when it is returned in time, and otherwise defaulting to the SVLM. This baseline represents the upper-bound performance of current cloud–edge collaborative VLM architectures. Compared to this baseline, we incrementally integrate each proposed module. Results show that CRM yields the most significant performance gain. GGVR is used to reduce the number of visual tokens, while GTVR guides the SVLM to maintain consistent detection of object categories present in the historical context. Building on CRM, the integration of these two methods further enhances overall system performance across all three metrics.
5.4 Adaptability Study
To evaluate the effectiveness of the proposed framework under different combinations of model capacities, we conducted additional experiments by replacing both the small and large models. As in the ablation study, all image frames are uploaded and processed by the large model, and the upper bound is represented by a parallel inference strategy that always adopts the output of large model when available. Specifically, we adopt MiniCPM-o (Yao et al. 2024) as the SVLM on the edge side and GPT-4o (Hurst et al. 2024) as the LVLM on the cloud side. As shown in the Tab.3, the proposed SpotVLM coordination strategy consistently improves real-time inference performance across model configurations. With the continuous advancement of small and large VLMs, more powerful models can be seamlessly integrated into the SpotVLM framework to further enhance system capabilities.
Figure 5: Latency simulation. Manual simulation of total processing latency on the cloud LVLM. Red pentagrams denote the performance bounds of LVLM and SVLM; the dashed line shows the SVLM-only performance.
5.5 Latency Analysis
To evaluate the impact of cloud latency on SpotVLM, we conduct experiments on the nuScenes dataset using Qwen-2.5-VL-3B as the edge-side SVLM and preserving the outputs of Qwen-2.5-VL-72B as the cloud-side LVLM. We manually simulate various latency conditions by setting different return delays for the LVLM responses.
When the latency is 1 second, which means that the LVLM output always arrives within the real-time threshold, the system achieves performance equivalent to directly using the LVLM (denoted by the star marker). Starting from a 2-second delay, the LVLM result can only be incorporated as historical context for guiding the SVLM. Even under this condition, SpotVLM significantly outperforms the standalone SVLM (indicated by the dashed line). As latency increases, the temporal correlation between frames weakens, leading to a gradual decline in overall performance. However, SpotVLM continues to maintain a notable advantage even under a 9-second delay.
This experiment reveals that the performance upper bound is achieved when the LVLM returns results in real time for every frame, while the lower bound corresponds to using only the SVLM. The proposed collaborative architecture offers strong robustness under fluctuating latency conditions. Moreover, the results also suggest that cloud-side LVLM invocation can be dynamically scheduled. By adjusting the upload interval according to task requirements and performance targets, the system can reduce the frequency of cloud model calls without significantly compromising overall effectiveness.
6 Conclusion
In this work, we propose a novel context transfer paradigm that enables LVLMs to provide real-time guidance to SVLMs through semantic and visual contexts. Building on this paradigm, we introduce SpotVLM, a collaborative framework that integrates two corresponding modules: Context Replacement Module and Visual Focus Module. Extensive experiments across diverse tasks, datasets, and model combinations demonstrate that SpotVLM outperforms existing collaboration methods. Moving forward, we envision integrating stronger foundation models and adaptive scheduling strategies into the collaborative framework to further optimize performance in real-world deployments.

### v2

@@
-SpotVLM: Cloud-edge Collaborative Real-time VLM based on Context Transfer
-SpotVLM: Cloud-edge Collaborative Real-time VLM
-based on Context Transfer
-Chen Qian1, Xinran Yu1, Zewen Huang1, Danyang Li1, Qiang Ma1,
-Fan Dang2, Xuan Ding1, Guangyong Shang3, Zheng Yang1
-1 Abstract
+edgeVLM: Cloud-edge Collaborative Real-time VLM based on Context Transfer
+Abstract
 Vision–Language Models (VLMs) are increasingly deployed in real-time applications such as autonomous driving and human–computer interaction, which demand fast and reliable responses based on accurate perception.
@@
 In this work, we propose a novel cloud–edge collaborative paradigm for VLMs, termed Context Transfer, which treats the delayed outputs of LVLMs as historical context to provide real-time guidance for SVLMs inference.
-Based on this paradigm, we design SpotVLM, which incorporates both context replacement and visual focus modules to refine historical textual input and enhance visual grounding consistency.
-Extensive experiments on three real-time vision tasks across four datasets demonstrate the effectiveness of the proposed framework.
+Based on this paradigm, we design edgeVLM, which incorporates both context replacement and visual focus modules to refine historical textual input and enhance visual grounding consistency.
+Extensive experiments on three real-time vision-language reasoning tasks across four datasets demonstrate the effectiveness of the proposed framework.
 The new paradigm lays the groundwork for more effective and latency-aware collaboration strategies in future VLM systems.
+Code will be publicly released before publication.
 1 Introduction
@@
 VLMs have witnessed rapid advancements in recent years.
-These models typically consist of a visual encoder, a language decoder, and a cross-modal fusion module, enabling them to perform a wide range of tasks such as image captioning (Xue et al.
-2025), visual question answering (Fang et al.
-2025), and robotic perception (Zhou et al.
-2025).
-Notably, LVLMs (Liu et al.
-2025), empowered by large-scale model capacity and extensive training data, demonstrate superior generalization ability and reasoning performance across a broad range of downstream tasks.
-However, due to their substantial computational demands, these models are often deployed on the cloud servers, which inevitably involve network transmission, model inference, and request queuing delays, posing challenges for scenarios that require real-time responses within seconds.
-To overcome these limitations, recent studies have proposed lightweight SVLMs for edge deployment (Yao et al.
-2024; Rang et al.
-2025), enabling low-latency, on-device inference suitable for real-time scenarios like autonomous driving (Lu et al.
-2025; Sima et al.
-2024; Tian et al.
-2024), augmented reality (Pan and Liu 2025), and smart wearable devices (Huang et al.
-2024).
-However, due to their limited model capacity, SVLMs still exhibit significantly lower inference accuracy than LVLMs (Zhao et al.
-2025).
-To enable efficient and accurate real-time visual reasoning, recent studies (Hu et al.
-2024b; Li et al.
-2025b; Hu et al.
-2024a) have explored cloud–edge collaborative architectures.
-As illustrated in Fig.1, existing approaches fall into two categories: (i) Model partitioning, which splits the LVLM across the cloud and edge based on hardware constraints.
-For example, Distributed VLM (Li et al.
-2025b) deploys the visual encoder on the edge and transmits extracted features to the cloud, where they are fused with text and processed by an LLM.
-(ii) Dynamic task offloading, which routes inputs based on task complexity and system conditions.
-For instance, ADAS (Hu et al.
-2024b) offloads tasks to the cloud for LVLM inference by considering factors such as latency and quality of service (QoS).
-Although such methods reduce cloud-side inference workload, they still face performance bottlenecks from service latency.
-In video streaming scenarios, accumulated delays often cause the outputs of LVLMs arriving several frames behind the current input (e.g., producing results for frame t–d while the system is processing frame t).
-Consequently, although these outputs are of high quality, they are temporally outdated and cannot be effectively utilized for real-time reasoning.
+These models typically consist of a visual encoder, a language decoder, and a cross-modal fusion module, enabling them to perform a wide range of tasks such as image captioning [xue2025progress], visual question answering [Fang_2025_CVPR], and robotic perception [zhou2025physvlm, liu2025reasongrounder].
+Notably, LVLMs [liu2025making], empowered by large-scale model capacity and extensive training data, demonstrate superior generalization ability and reasoning performance across a broad range of downstream tasks.
+However, due to their substantial computational demands, these models are often deployed on the cloud servers, as illustrated in Fig.1(b).
+This deployment inevitably involve network transmission, model inference, and request queuing delays, posing challenges for scenarios that require real-time responses within seconds.
+To overcome these limitations, recent studies have proposed lightweight SVLMs for edge deployment [yao2024minicpm, rang2025eve, lu2025bluelm], enabling low-latency [vasu2025fastvlm], on-device inference suitable for real-time scenarios like autonomous driving [lu2025can, sima2024drivelm, tian2024drivevlm], augmented reality [pan2025metaspatial], and smart wearable devices [huang2024vinci], as illustrated in Fig.1(a).
+However, due to their limited model capacity, SVLMs still exhibit significantly lower inference accuracy than LVLMs [zhao2025stitch].
+To enable efficient and accurate real-time vision-language reasoning, recent studies [hu2024cloud, li2025distributed, hu2024laecips] have explored cloud–edge collaborative architectures.
+Existing approaches fall into two categories: (i) Model partitioning , which splits the LVLM across the cloud and edge based on hardware constraints.
+A representative example is Distributed VLM [li2025distributed], depicted in Fig.1(c), which deploys the visual encoder on the edge and transmits extracted features to the cloud.
+(ii) Dynamic task offloading, which routes inputs based on task complexity and system conditions, as illustrated in Fig.1(d).
+For instance, ADAS [hu2024cloud] offloads tasks to the cloud for LVLM inference by considering factors such as latency and quality of service (QoS).
+However, these approaches inherently depend on obtaining timely feedback from the LVLM to produce real-time outputs for the current frame, and they largely ignore the additional latency caused by network jitter and request queuing in cloud services.
 To further explore more effective collaboration paradigms between large and small VLMs, we analyze the underlying factors that contribute to their performance gap and derive the following insights:
@@
 Historical context plays a crucial role in generating high-quality outputs.
-For example, Qwen-VL (Bai et al.
-2023) adopts an accumulative input structure in multi-turn interactions: each round of input consists of the input and response from the previous round, along with the current query.
+Qwen-VL [bai2023qwen], as a representative VLM, adopts an accumulative input structure in multi-turn interactions: each round of input consists of the input and response from the previous round, along with the current query.
 This context is progressively extended until the window reaches the maximum memory or model-defined limit.
 Such accumulation allows the model to integrate information across turns, enhancing semantic consistency and reasoning depth.
-More importantly, the generated outputs are recursively fed into subsequent inputs, forming a performance-enhancing feedback loop.
-Although small models adopt similar input strategies, their limited generative capacity often results in lower-quality historical content, which negatively affects reasoning performance.
+Due to its limited generative capacity, SVLM often produces low-quality or even erroneous historical content, which propagates to subsequent steps and degrades its reasoning performance.
 Enhancing the quality of contextual history thus emerges as a key factor in improving the reasoning ability of small models.
 $\bullet$ Grounding of Semantically Salient Image Regions.
-The ability to align visual content with textual descriptions is critical in vision-language reasoning.
+The ability to align visual content with textual descriptions is critical in vision-language reasoning [zong2025ground].
 However, SVLMs often struggle in this aspect, particularly when dealing with fine-grained semantics or complex scenes.
-Prior studies (Zheng et al.
-2025; Xu et al.
-2025; Zhang et al.
-2025a) have demonstrated that LVLMs, leveraging their grounding capabilities, can automatically crop semantically rich image regions as refined inputs—thereby enhancing the focus and semantic interpretability in visual reasoning.
+Prior studies [zheng2025deepeyes, xu2025visual, zhang2025chain] have demonstrated that LVLMs, leveraging their grounding capabilities, can automatically crop semantically rich image regions as refined inputs, thereby enhancing the focus and semantic interpretability in visual reasoning.
 Guiding small models to focus on task-relevant image regions enhances their ability to interpret complex visual and linguistic cues, ultimately improving their reasoning performance.
-Our Work: Given the inherent limitations of cloud-based LVLMs in terms of latency and response stability, we propose a key design principle in this work: Context Transfer.
-Instead of relying on the delayed outputs from LVLMs as real-time feedback, they are reused as contextual priors to enhance future reasoning in SVLMs.
-Building on this principle, we introduce SpotVLM, which incorporates two collaborative modules that guide the SVLM along both the linguistic and visual dimensions to achieve more accurate multimodal reasoning.
+Our Work: Motivated by the critical role of accurate and task-relevant context in model reasoning, we propose a key design principle in this work: Context Transfer.
+Instead of relying on the delayed outputs from LVLMs as real-time feedback, they are reused as contextual priors to enhance future reasoning in SVLMs, as illustrated in Fig.1(e).
+Building on this principle, we introduce edgeVLM, which incorporates two collaborative modules that guide the SVLM along both the linguistic and visual dimensions to achieve more accurate multimodal reasoning.
 Motivated by the strong impact of historical context quality on VLM inference, we design a Context Replacement Module (CRM).
 Specifically, It extracts key textual segments from high-quality LVLM responses, reorganizes them based on task context, and embeds them as contextual history to replace the original low-quality history.
-These informative exemplars implicitly guide the small model to focus on semantically relevant visual cues that were previously overlooked, thereby enhancing cross-modal association and reasoning performance.
+By exploiting inter-frame correlations among temporally adjacent frames, these refined historical cues offer precise contextual guidance for current-frame inference, implicitly steering the small model toward semantically relevant visual evidence that was previously overlooked.
 To exploit the grounding capabilities of large models, we introduce the Visual Focus Module (VFM), which identifies semantically salient regions from historical images based on LVLM predictions.
@@
 Semantically aligned regions are assigned higher weights, allowing the SVLM to inherit and shift attention toward task-relevant areas in the current frame.
-By providing high-quality textual history and guiding attention to semantically relevant visual regions, these modules effectively leverage the delayed outputs of LVLMs to compensate for the limited reasoning capacity of SVLMs.
 The contributions of this work are summarized as follows:
-  • We propose a novel architectural design paradigm, termed Context Transfer, for real-time vision-language reasoning.
+  • We propose Context Transfer, a cloud–edge collaborative paradigm for real-time VLM inference.
 Instead of relying on the final predictions from LVLMs, this paradigm treats their delayed outputs as contextual priors to guide SVLMs, enabling a new form of efficient large–small model collaboration.
-  • Based on this principle, we design SpotVLM with two zero-shot collaborative modules: the Context Replacement Module and the Visual Focus Module, which guide the SVLMs to reuse high-quality outputs from the LVLMs along the linguistic and visual dimensions, respectively.
-  • Extensive experiments on four public benchmarks validate the effectiveness of SpotVLM, showing consistent improvements over prior cloud–edge collaborative approaches.
-3 Related Work
-3.1 Cloud-Edge Collaborative VLMs
-Figure 2: Overview of the proposed collaboration framework, SpotVLM.
+  • Based on this principle, we design edgeVLM with two training-free collaborative modules: the Context Replacement Module and the Visual Focus Module, which guide the SVLMs to reuse high-quality outputs from the LVLMs along the linguistic and visual dimensions, respectively.
+  • Extensive experiments on four public benchmarks validate the effectiveness of edgeVLM, showing consistent improvements over prior cloud–edge collaborative approaches.
+Figure 2: Overview of the proposed collaboration framework, edgeVLM.
 The system takes timestamped video streams as input and performs two parallel operations: uploading selected keyframes to the cloud-based LVLM for processing, and conducting local inference using the SVLM.
@@
 These outputs guide the SVLM through two modules, Context Replacement and Visual Focus, to improve the quality of real-time predictions.
-Driven by the growing demand for real-time VLM-based processing, applications such as autonomous driving (Lu et al.
-2025) are shifting from modular to end-to-end pipelines (Shao et al.
-2024; Li et al.
-2025a), where latency sensitivity is critical and excessive delays can lead to system failure.
+2 Related Work
+2.1 Cloud-Edge Collaborative VLMs
+Driven by the growing demand for real-time VLM-based processing, applications such as autonomous driving [lu2025can] are shifting from modular to end-to-end pipelines [shao2024lmdrive, li2025generative], where latency sensitivity is critical and excessive delays can lead to system failure.
 To enable fast and robust responses, cloud-edge collaborative architectures for VLMs have attracted increasing attention.
 Existing collaboration paradigms can be categorized into two types.
-The first category is model partitioning (Li et al.
-2025b; Zhang et al.
-2025b).
-To improve processing efficiency for LVLMs in the cloud, Distributed VLM (Li et al.
-2025b) offload the visual encoder to the edge device, which encodes input images rapidly and transmits the feature embeddings to the cloud for processing and decoding.
-VaVLM (Zhang et al.
-2025b) further reduces transmission overhead by uploading only the region-of-interest (ROI) of images.
-This parallel architecture entirely rely on stable and responsive cloud services, making it difficult to guarantee robust frame-wise inference in real-time settings.
+The first category is model partitioning [li2025distributed, zhang2025vavlm].
+To improve processing efficiency, Distributed VLM [li2025distributed] offload the visual encoder to the edge device, which encodes input images and transmits features to the cloud for processing and decoding.
+VaVLM [zhang2025vavlm] reduces transmission overhead by uploading only the region-of-interest (ROI) of images.
+These parallel architecture entirely rely on stable and responsive cloud services, making it difficult to guarantee robust frame-wise inference in real-time settings.
 The second category involves dynamic task allocation to different models according to predefined offloading strategies.
-Some systems (Hu et al.
-2024b) formulate the offloading decision as an optimization problem, incorporating multiple system-level factors such as Quality of Service (QoS) and latency.
-In addition, frameworks such as LAECIPS (Hu et al.
-2024a) pretrain a dedicated task classifier to estimate task complexity and offload difficult samples to cloud-based large models.
+Some systems [hu2024cloud] formulate the offloading decision as an optimization problem, incorporating multiple system-level factors such as Quality of Service (QoS) and latency.
+Frameworks such as LAECIPS [hu2024laecips] pretrain a dedicated task classifier to estimate task complexity and offload difficult samples to cloud-based large models.
 While these approaches leverage the complementary strengths of large and small models to reduce cloud-side computation, they overlook the continued guidance that high-quality LVLM responses can provide for subsequent inference.
@@
 In VLMs, this mechanism is adapted from Large Language Models (LLMs), which use fixed-length context windows to preserve relevant token sequences and enhance reasoning consistency.
-Prior research on LLMs has shown that adjusting context improves model performance, as reflected in translation accuracy (Sung et al.
-2024) and dialogue consistency (Wei et al.
-2024).
-Similarly, VLMs leverage effective context preservation to significantly improve robotic navigation in embodied AI scenarios (Habibpour and Afghah 2025), reducing the risk of decision paralysis and oscillatory behavior.
-Building on this advantage, the proposed SpotVLM retain the output from LVLMs as high-quality context to guide the SVLMs, helping it focus on salient content and improve inference quality.
+Prior research on LLMs has shown that adjusting context improves model performance, as reflected in translation accuracy [sung2024context] and dialogue consistency [wei2024hidden].
+Similarly, VLMs leverage effective context preservation to significantly improve robotic navigation in embodied AI scenarios [habibpour2025history], reducing the risk of decision paralysis and oscillatory behavior.
+Building on this advantage, the proposed edgeVLM retain the output from LVLMs as high-quality context to guide the SVLMs, helping it focus on salient content and improve inference quality.
 3 Architecture
 Real-time vision-language reasoning tasks pose dual challenges in terms of processing speed and inference accuracy.
-To achieve robust and high-precision outputs, this work leverages the strengths of LVLMs in textual understanding and visual perception to assist SVLMs in real time.
-This section illustrates the proposed collaborative architecture using a multi-object recognition task in autonomous driving at a frame rate of 1 FPS as a representative example.
+To achieve robust and high-precision outputs, this work leverages the strengths of LVLMs in textual understanding and visual perception to assist SVLMs in real-time.
+This section illustrates the proposed collaborative architecture using a multi-object recognition task in autonomous driving as a representative example.
 3.1 Workflow
 The front-facing camera continuously captures frames and sends them to the edge device.
-For each timestamped frame, the edge decides whether to upload it to the cloud based on the latency of the previous cloud interaction.
+For each timestamped frame, the edge decides whether to upload it to the cloud.
 Each frame is also processed locally by the SVLM for real-time inference.
-For uploaded frames, LVLM responses are handled according to latency: results within 1 second are adopted and stored, while delayed outputs are retained for later use, and real-time results from the SVLM are adopted for the current frame.
-As illustrated in Fig.2, the output returned by the LVLM contains two key components: (1) the response content to the given query, denoted as <answer>; and (2) the attended image regions during inference, indicated as <ROI>.
-These two outputs provide guidance for subsequent frames from textual and visual perspectives, respectively.
-In addition, the returned result includes the timestamp of the original frame, which allows the system to assess current cloud latency and determine the temporal relevance of the result for semantic guidance.
+For uploaded frames, the returned LVLM responses are handled according to latency: results within $\tau$ second are adopted and stored, while delayed outputs are retained for later use, and real-time results from the SVLM are adopted for the current frame.
+As illustrated in Fig.2, the output for frame TdT-d returned by the LVLM contains two key components: (1) the response content to the given query, denoted as <answer>; and (2) the attended image regions during inference, indicated as <ROI>.
+The two outputs respectively provide textual and visual guidance for SVLM inference at frame $T$.
 3.2 Context Replacement
@@
 To support continuous multimodal understanding, VLMs often retain part of the historical context during inference.
-This insight informs the design of two novel mechanisms for enabling textual-level collaboration across LVLMs and SVLMs.
-As detailed in Sec.4.1, high‑quality outputs from the large model are continuously updated and stored in an edge‑side cache as $\mathcal{H}_{T-d}^{LVLM}$:
-$\mathcal{H}_{T-d}^{LVLM}\leftarrow LVLM(I_{T-d},Q_{T-d}),$ (1)
-where $I_{T-d}$ and $Q_{T-d}$ denote the image and question for Frame T-d, respectively.
-When processing frame T, the system compares the current timestamp with the most recent timestamp T-d associated with a returned LVLM result.
-If the delay is within a predefined threshold $\delta$, the SVLM replaces its original output for frame T-d with the response from LVLM, and incorporates it into the historical context to guide inference on frame T.
-$if\quad\Delta t=T-(T-d)\leq\delta,\quad\mathcal{H}_{T-d}^{SVLM}\leftarrow\mathcal{H}_{T-d}^{LVLM}$ (2)
-Given the temporal continuity and strong correlation between adjacent frames, providing more accurate historical information increases the likelihood of correctly activating relevant tokens and suppressing irrelevant ones, thereby improving the overall precision of outputs generated by the SVLM.
-Moreover, instead of directly concatenating raw questions and answers as historical context, we design an in-context learning prompt template to compress and summarize dialogue history for the SVLM, and the result is denoted as follows:
-$\tilde{\mathcal{H}}=Compress(\mathcal{H}_{\leq T-1}).$ (3)
+We find that replacing the history generated by SVLM with accurate historical context supplied by a delayed LVLM guides the SVLM to produce more accurate predictions for the current frame.
+To theoretically justify this empirical observation, consider the following sequential prediction formulation.
+Let the prediction at time $t$ be
+$\hat{\mathbf{y}}_{t}=f(\mathbf{x}_{t},\mathbf{h}_{t-1}),$ (1)
+where $\mathbf{x}_{t}$ is the current input and $\mathbf{h}_{t-1}$ is the historical state from frame (t1)(t\!-\!1).
+Denote by $\mathbf{h}^{\star}_{t-1}$ the accurate history and by $\mathbf{h}^{\star}_{t-1}+\mathbf{e}$ an inexact history.
+A first-order expansion in the historical state gives
+$f(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1}+\mathbf{e})\approx f(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1})+J_{h}\,\mathbf{e},$ (2)
+$J_{h}:=\frac{\partial f}{\partial\mathbf{h}}\Big|_{(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1})}.$ (3)
+Hence the induced output deviation is bounded by
+$\big\|f(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1}+\mathbf{e})-f(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1})\big\|\leq\|J_{h}\|\,\|\mathbf{e}\|.$ (4)
+Let $\ell=\big\|\cdot\big\|^{2}$, the expected excess loss at time $t$ satisfies
+$\mathbb{E}[\Delta\ell]\;\lesssim\;\mathrm{tr}\!\big(J_{h}\,\Sigma_{e}\,J_{h}^{\top}\big),$ (5)
+where $\Sigma_{e}$ is the covariance matrix of the historical error.
+Therefore, a more accurate history (smaller $\Sigma_{e}$) yields a smaller expected loss at frame $t$.
+This insight informs the design of two novel methods for enabling textual-level collaboration across LVLMs and SVLMs.
+As detailed in Sec.
+3.1, the <answer> produced by the LVLM is continuously refreshed and preserved in the edge-side cache.
+At frame T, the system locates in the edge cache the latest LVLM result corresponding to a prior frame T - d.
+If the elapsed time $d\leq\delta$, that LVLM result replaces the historical output produced by the SVLM for frame T - d and is incorporated into the context used to condition inference on frame T.
+Since adjacent frames are highly correlated, higher-quality historical information more reliably steers the model toward activating relevant tokens and suppressing irrelevant ones.
 Given the limited memory capacity of edge devices and the context window constraints of lightweight models, naively stacking multi-turn inputs may lead to token truncation.
-By compressing the historical content, our method not only reduces the number of irrelevant tokens but also enhances attention to the semantics of prior answers for the model.
-To summarize, with the textual collaboration strategy in place, the SVLM receives an input structure at frame T that corresponds to the right side of the Context Replacement module in the figure.
-$Input_{T}^{SVLM}=\{I_{T},Q_{T},\tilde{\mathcal{H}}\}$ (4)
-This input incorporates distilled historical answers from the LVLM, enabling the SVLM to perform accurate and efficient semantic reasoning.
-Figure 3: ROI-based region tracking via cosine similarity.
-Given ROIs of traffic cone and pedestrian from frame T–d, cosine similarity highlights related regions in frame T to guide the attention of the small model.
+Moreover, we design an in-context learning [kang2025context] prompt template to compress and summarize dialogue history for the SVLM.
+Example.
+In the autonomous driving setup, the edge-side SVLM predicts “car” on the historical frame T - d.
+If the corresponding cloud response arrives within the staleness bound $d\leq\delta$, we replace the history with the corrected set from LVLM “traffic cone, car, pedestrian.” We then convert the Q&A-style history into a compressed snippet: “This is the image from d seconds ago; detected traffic cone, car, pedestrian.” This snippet is injected into the current input as prior.
+To summarize, with the textual collaboration strategy in place, the SVLM receives an input structure at frame T that corresponds to the right side of the Context Replacement module in Fig.
+2.
+This input incorporates distilled historical answers from the LVLM, enabling the SVLM to perform accurate and efficient reasoning.
 3.3 Visual Focus
@@
 Based on this, SVLMs can proactively focus on critical visual regions from two perspectives.
-Grounding-Guided Visual Representation.
+Figure 3: ROI-based region retrieving via cosine similarity.
+Given ROIs of traffic cone and pedestrian from frame T – d, cosine similarity highlights related regions in frame T to guide the attention of the SVLM.
+3.3.1 Grounding-Guided Visual Representation
 The ROIs returned by the LVLM guide the grounding of the SVLM on historical frames.
-Most VLMs, including the Qwen-2.5-VL-3B (Bai et al.
-2025), use a Vision Transformer (ViT) that divides a resized input image into patches and encodes them into visual tokens (e.g., 1824 tokens for a 1600 × 900 image).
-However, many of these tokens are irrelevant to the reasoning task.
+Most VLMs, including the Qwen-2.5-VL-3B [Qwen2.5-VL], use a Vision Transformer (ViT) [dosovitskiy2020image] that divides a resized input image into patches and encodes them into visual features (e.g., 1824 features for a 1600 × 900 image).
+However, many of these features are irrelevant to the reasoning task.
 To accelerate inference and enhance focus, ROIs are mapped to corresponding patches, and only task-relevant ones are selected.
@@
 In the case of multi-object recognition, the largest ROI is selected for each target category.
-For the case illustrated in the figure, the number of features selected for projection is reduced to 49, thereby reducing the computational cost.
-nuScenes BDD100K MOT IPN Hand ACE
-Micro-F1 Macro-F1 0-1 EM Micro-F1 Macro-F1 0-1 EM Acc. Acc.
-SVLM Only (Local) 58.19 37.70 16.10 62.12 31.06 6.37 36.41 23.43
-LVLM Only (Cloud) 45.76 32.04 20.75 49.65 25.85 20.28 23.91 61.14
-Distributed VLM (Li et al. 2025b) 45.31 31.78 20.39 49.31 25.71 20.09 23.64 60.86
-LAECIPS (Hu et al. 2024a) 59.84 39.64 24.29 62.47 32.14 21.05 39.40 48.71
-ADAS (Hu et al. 2024b) 60.00 41.35 18.11 63.87 33.12 10.63 39.27 55.43
-SpotVLM 69.96 48.19 27.32 71.19 36.26 23.88 45.11 62.75
-Table 1: Performance comparison of different architectures across four datasets (values in %). All methods adopt Qwen-VL-2.5-3B as the SVLM and Qwen-VL-2.5-72B as the LVLM.
-Grounding-Tracked Visual Representation.
-The ROIs provided by the LVLM support cross-frame tracking, guiding the SVLM to focus on relevant regions in subsequent frames.
-Due to limited grounding capability, lightweight SVLMs may struggle to localize objects even when their categories are known.
-However, given the temporal continuity of video, ROI features from frame T-d can be matched to visual features in frame T.
+For the case illustrated in Fig.
+2, the number of visual tokens selected after projection is reduced to 49, thereby reducing the computational cost.
+3.3.2 Grounding-Retrieved Visual Representation
+The ROIs provided by the LVLM support cross-frame retrieving, guiding the SVLM to focus on relevant regions in subsequent frames.
+Given the temporal continuity, ROI features from frame T - d can be matched to visual features in frame T.
 Let $\mathcal{P}_{T-d}^{(c)}$ denotes the set of patch indices for category $c$ on frame T - d, and $\mathbf{v}_{i}^{(T-d)}$ is the visual feature of the $i$-th patch.
@@
 $\mathbf{f}^{(c)}=\frac{1}{|\mathcal{P}^{(c)}_{T-d}|}\sum_{i\in\mathcal{P}^{(c)}_{T-d}}\mathbf{v}^{(T-d)}_{i}$. (7)
-We compute the cosine similarity between ROI features from frame T-d and the visual features of frame T, thereby obtaining similarity maps $S^{(c)}$ for each potential category $c$ in frame T.
+We compute the cosine similarity between ROI features from frame T - d and all visual features from frame T, thereby obtaining similarity maps $S^{(c)}$ for each potential category $c$ in frame T.
 $\mathbf{S}^{(c)}=\left[\frac{\mathbf{f}^{(c)}\cdot\mathbf{v}^{(T)}_{j}}{\|\mathbf{f}^{(c)}\|\cdot\|\mathbf{v}^{(T)}_{j}\|}\right]_{j=1}^{N},$ (8)
@@
 As shown in Fig.3, the LVLM accurately detects and localizes small objects such as traffic cones and pedestrians in frame T - d, and their mapped similarity heatmaps on frame T effectively indicate the potential locations after viewpoint shift, providing precise visual guidance for the SVLM.
-This cross-frame computation yields patch-level similarity scores can be aggregated into a unified similarity vector $s$, which can be defined as:
+This cross-frame computation yields patch-level similarity scores, which can be aggregated into a unified similarity vector $s$, defined as:
 $s=\frac{1}{C}\sum_{c=1}^{C}S^{(c)}.$ (9)
-Then $s$ can be used to construct the grounding-tracked visual weight $W_{s}$, formulated as follows:
+Then $s$ can be used to construct the grounding-retrieved saliency weight $W_{s}$, formulated as follows:
 $W_{s}=\mu\cdot\frac{1}{1+e^{-s}}+b,$ (10)
-where $\mu$ and $b$ are tunable scaling parameters used to control the visual weights.
+where $\mu$ and $b$ are tunable scaling parameters used to control the visual weights. Here, we enforce $0.5\mu+b=1$ in our experiments.
 To further adapt the visual representation based on semantic associations, we introduce a gated token update mechanism that enables dynamic modulation of visual tokens. The update is formulated as:
-$t\leftarrow(1-\alpha)\cdot t+\alpha\cdot W_{s}\cdot t.$ (10)
+$t\leftarrow\alpha\cdot t+(1-\alpha)\cdot W_{s}\cdot t.$ (11)
 Here, $t$ denotes the original visual token, and $\alpha\in[0,1]$ is a control coefficient that balances the weights between known object categories and newly emerging targets.
 The dynamically predicted weight $W_{s}$ modulates the visual tokens based on feature similarity, enabling adaptive token refinement under the guidance of semantic consistency.
+3.4 Keyframe Decision
+Figure 4: Experimental latency evaluation of cloud services over a 5G network.
+Under a 5G network environment, we invoke the Qwen-VL-2.5-72B API hosted on Alibaba Cloud and record the per-frame response latency, as illustrated in the Fig.4.
+The results show that adjacent frames exhibit similar latency patterns.
+To enhance the efficiency of LVLM invocations and dynamically assess network conditions, the system continuously monitors how many preceding frames have not yet received responses.
+Based on this observation, we introduce a hyperparameter $K_{delay}$, which serves as a threshold: when the number of unreturned frames accumulates to $K_{delay}$, the system infers that the network condition has deteriorated and skips uploading the current frame.
+We further regulate LVLM call frequency with a hyperparameter $K_{call}$.
+For frame $T$, if the most recent LVLM result is from frame TiT-i with $i\leq K_{call}$, that history is considered sufficiently fresh to guide the SVLM, and frame $T$ is not uploaded.
+This rule takes precedence over $K_{delay}$.
 4 Experiments
 4.1 Experiments Setting
-Datasets.
-We evaluate the proposed architecture on three real-time tasks.
+4.1.1 Datasets
+We evaluate the proposed architecture on three real-time tasks across four datasets.
 Real-time multi-object recognition.
-(i) nuScenes (Caesar et al.
-2019): We sample camera images at 1 FPS from this autonomous driving dataset.
-The processed data includes 150 scenes (20 seconds each) with 10 object categories.
-(ii) BDD100K (MOT) (Yu et al.
-2020): We sample the driving video frame at 1 FPS.
-The processed data includes 200 scenes (40 seconds each) with 10 object categories.
+(i) nuScenes [nuscenes2019]: We sample images at 1 FPS from this autonomous driving dataset, which finally includes 156 scenes (avg.
+20s each) with 10 object categories.
+(ii) BDD100K (MOT) [yu2020bdd100k]: We sample the driving video frame at 1 FPS, which includes 200 scenes (avg.
+40s each) with 10 object categories.
 Real-time gesture recognition.
-IPN Hands (Benitez-Garcia et al.
-2021): A gesture recognition dataset with 14 gesture classes, sampled at 2 FPS from 52 close-range videos (avg.
+IPN Hands [bega2020IPNhand]: A gesture recognition dataset with 14 gesture classes, sampled at 2 FPS from 52 close-range videos (avg.
 120s each) to capture dynamic motion.
-Real-time video frame caption. Actions for Cooking Eggs (ACE) (Shimada et al. 2012): A kitchen scene dataset with 8 cooking actions, sampled at 1 FPS from 10 videos (avg. 194s each).
-Metrics.
-We adopt commonly used evaluation metrics (Durand, Mehrasa, and Mori 2019; Nam et al.
-2017; Huang et al.
-2017) for multi-object recognition task, which are Micro-F1, Macro-F1 and 0-1 Exact Match (0-1 EM).
-Specifically, Micro-F1 reflects overall performance across all instances, and Macro-F1 captures class-wise performance equally.
-Additionally, 0-1 EM metric equals 1 only if the predicted result exactly matches the ground truth label for the current frame.
-Following 0-1 EM, we model both gesture recognition and video frame captioning as frame-level single-label classification tasks, and use accuracy (Acc) as the evaluation metric.
-Note that the evaluated tasks operate under real-time constraints (1–2 FPS). If the system fails to return a valid result within one second, the output for that frame is considered empty.
+Real-time video frame caption. Actions for Cooking Eggs (ACE) [shimada2012kitchen]: A kitchen scene dataset with 8 cooking actions, sampled at 1 FPS from 10 videos (avg. 194s each).
+4.1.2 Metrics
+We adopt commonly used evaluation metrics [durand2019learning, nam2017maximizing, huang2017multi] for multi-object recognition task, which are Micro-F1, Macro-F1 and 0-1 Exact Match (0-1 EM).
+0-1 EM equals 1 only if the predicted result exactly matches the ground truth label.
+We model both gesture recognition and video frame captioning as frame-level single-label classification tasks, and use accuracy (Acc) as the evaluation metric.
+Note that the evaluated tasks operate under real-time constraints (1–2 FPS). Following DriveVLM [tian2024drivevlm], we set $\tau=1$ in experiments.
+4.1.3 Implementation Details
+Experiments are conducted over a commercial 5G network.
+The average uplink bandwidth was 211 Mbps, downlink 72 Mbps, with measured round-trip latency around 42 ms.
+To remove variability from network jitter and queuing, we execute the cloud-side LVLM on every frame and log its streaming outputs and the per-frame latency, as shown in Fig.
+4.
+All collaboration strategies are evaluated against these identical per-frame traces for a fair comparison.
+To balance processing speed and model performance, we retain two previous frames as context. All experiments on the edge side are conducted using a GeForce RTX 4090 GPU with 24 GB of memory.
+The following abbreviations are used in this section: Grounding-Guided Visual Representation (GGVR), and Grounding-Retrieved Visual Representation (GRVR).
+4.2 Hyperparameter Study
+LVLM Call.
+To balance the system performance and the cost of invoking the LVLM, we introduce two hyperparameters, $K_{delay}$ and $K_{call}$, in the Keyframe Decision module.
+A larger $K_{delay}$ leads to more frequent LVLM invocations.
+We conduct evaluations on the nuScenes dataset, and the resulting system performance is shown in Fig.5(a).
+$K_{delay}=4$ is selected as a trade-off choice.
+Under the 5G network, we recorded the LVLM utilization under different $K_{call}$ settings, as shown in Tab.1.
+LCR (LVLM Call Ratio) is the fraction of inference steps that trigger the cloud LVLM.
+LHL (LVLM History Lag) is the average frame offset $i$ between the current frame $t$ and the reused LVLM reply from frame tit-i, computed over all frames that actually consume LVLM history.
+When $K_{call}$ increases from 2 to 3, the LVLM invocation frequency drops sharply, while the overall system performance remains relatively stable as shown in Fig.5(b).
+Visual Guidance. In the GRVR module, $\alpha$ and $\mu$ are introduced to regulate the visual guidance. Considering multi-metric performance and robustness, we adopt $(\alpha,\mu)=(0.7,1.0)$.
+Figure 5: Absolute performance gain of edgeVLM over SVLM across different settings. (a) shows the impact of $K_{delay}$ with $K_{call}=0$. (b) represents the impact of $K_{call}$ with $K_{delay}=4$.
+$K_{call}$ 1 2 3 4 5 6 7 8 9
+LCR(%) 69.1 38.9 27.3 23.5 19.0 18.4 17.8 14.4 13.0
+LHL(s) 1.89 2.56 3.31 3.79 4.12 4.54 5.14 5.48 5.66
+Table 1: Impact of $K_{call}$ on LVLM utilization.
+Method nuScenes BDD100K MOT IPN Hand ACE
+Micro-F1 Macro-F1 0-1 EM LCR Micro-F1 Macro-F1 0-1 EM LCR Acc LCR Acc LCR
+SVLM Only (Edge) 55.36 38.51 13.80 0 61.56 30.79 6.59 0 36.68 0 22.86 0
+LVLM Only (Cloud) 43.84 33.81 18.98 100 49.79 24.59 21.77 100 22.55 100 60.14 100
+Distributed VLM [li2025distributed] 43.35 33.63 18.84 100 48.28 23.55 20.89 100 21.47 100 59.43 100
+LVLM Call Ratio (20-30%)
+LAECIPS [hu2024laecips] 58.06 39.01 17.04 29.3 61.70 30.48 12.88 28.7 38.86 28.8 32.57 28.0
+ADAS [hu2024cloud] 56.43 40.28 18.43 28.2 62.10 31.54 13.61 29.5 37.09 25.6 33.29 28.5
+edgeVLM 64.49 48.00 21.99 27.3 68.44 34.76 23.66 28.4 46.20 20.7 34.71 25.9
+LVLM Call Ratio (35-45%)
+LAECIPS 57.24 39.39 19.58 44.5 61.88 30.90 15.98 44.6 37.23 44.0 40.29 44.7
+ADAS 57.52 41.13 20.03 43.7 62.55 30.62 16.88 44.2 35.87 43.8 39.71 43.6
+edgeVLM 66.22 50.04 24.51 43.2 70.63 36.41 26.88 43.7 47.42 35.4 38.86 35.1
+LVLM Call Ratio (50-60%)
+LAECIPS 56.28 39.06 21.18 55.3 61.91 30.28 17.56 59.2 34.51 58.7 45.71 54.0
+ADAS 58.03 42.23 22.13 55.9 62.88 30.62 19.99 57.4 35.33 56.2 43.23 54.8
+edgeVLM 67.24 51.35 25.00 52.8 71.70 37.31 29.16 57.3 49.73 55.7 46.14 52.9
+LVLM Call Ratio (100%)
+edgeVLM 71.15 57.11 29.41 100 74.12 39.66 35.08 100 52.45 100 63.00 100
+Table 2: Performance comparison of different architectures across four datasets (values in %). All methods adopt Qwen-VL-2.5-3B as the SVLM and Qwen-VL-2.5-72B as the LVLM.
 Figure 6: Qualitative Results. Illustration of input questions, ground-truth answers, and system predictions from four datasets.
-Implementation Details.
-In the Visual Focus Module, the parameters $\mu$, $b$ and $\alpha$ are set to 0.5, 0.75, and 0.1, respectively.
-To balance processing speed and model performance, we retain two previous frames as context.
-SpotVLM consistently treats the cloud feedback as the farthest frame in the historical context of SVLM.
-In addition, to evaluate the effectiveness of different collaborative architectures, we use the official pretrained versions of both the LVLMs and SVLMs without any additional fine-tuning.
-During inference, the outputs and response latencies of the large model are recorded for each frame to ensure fair comparison of collaboration strategies.
-All experiments on the edge side are conducted using a GeForce RTX 4090 GPU with 24 GB of memory.
-CRM GGVR GTVR Micro-F1 Macro-F1 0-1 EM
-66.87 43.07 24.39
-71.46 48.64 29.98
-72.15 50.85 30.93
-73.02 51.56 33.17
+4.3 System-level Evaluation
+4.3.1 Overall Performance
+To evaluate the effectiveness of different architectures, we assess six methods: edge-based Qwen2.5‑VL‑3B, cloud-based Qwen2.5‑VL‑72B [Qwen2.5-VL], Distributed VLM, LAECIPS, ADAS, and edgeVLM.
+The results are presented in Tab.2.
+For the fourth task, single-image inputs and short outputs allow rapid LVLM feedback, making full LVLM invocation clearly advantageous.
+For the other tasks, due to cloud latency, methods that rely entirely on LVLM yield lower F1 scores than collaborative approaches.
+However, the 0–1 EM of LVLM Only remains higher than that of SVLM Only, showing its superior accuracy when responses are timely.
+Additionally, Qwen2.5-VL-3B produces visual features larger than the original image, making Distributed VLM less effective than directly invoking the LVLM.
+Figure 7: Adaptability Study.
+Testing the edgeVLM architecture with various SVLM–LVLM combinations on nuScenes.
+VLM-R1 is a small model further enhanced by reinforcement learning and knowledge distillation.
+By tuning the hyperparameters, we evaluate the system performance of edgeVLM under different LCR settings and compare it with alternative architectures.
+ADAS adopts latency-adaptive offloading strategies, where LVLM invocation depends on the current cloud service state.
+As a result, it tends to use a relatively low LCR on the first three datasets under realistic deployment.
+In contrast, LAECIPS triggers LVLM calls based on task difficulty and thus chooses a higher LCR on these datasets.
+Overall, however, both schemes deliver lower practical performance than edgeVLM, which more effectively exploits historical information even with a comparatively low LCR.
+To provide a more intuitive comparison, we select four representative examples for qualitative analysis.
+As shown in Fig.6, for the real-time multi-object recognition task, all three methods rely on the SVLM, with edgeVLM achieving the best performance due to guidance from the historical results of the LVLM.
+In the gesture recognition task, as two images need to be transmitted and processed simultaneously, the LVLM fails to return valid results in real-time, leading to No response for ADAS.
+For the video-frame captioning task, both ADAS and edgeVLM adopt the results quickly returned by the LVLM.
+4.3.2 Latency Analysis
+System Stage Time (ms)
+LVLM Only Total 2110.83
+SVLM Only Total 581.71
+edgeVLM CRM-induced Overhead 19.93
+VFM-induced Overhead 31.27
+Inference 511.20
+Total 562.40
+Table 3: Execution time breakdown by system and stage (ms).
+To assess module-induced computational overhead, we report the average execution time of different systems in Tab.
+3.
+CRM and VFM introduce additional processing latency.
+However, by reducing the number of visual tokens, VFM lowers the compute over the entire inference pipeline and offsets part of this added overhead.
+As a result, edgeVLM attains a lower total latency than running the SVLM.
+Moreover, efficiency details are presented in Tab.4.
+System GPU Memory (MiB) Prefill Time (ms) Time / Token (ms)
+SVLM Only 10411 169.38 46.86
+edgeVLM 9589 166.81 44.75
+Table 4: Efficiency analysis of edgeVLM.
+4.4 Method-level Evaluation
+4.4.1 Ablation Study
+We conduct a progressive ablation study on the nuScenes dataset to evaluate the individual and combined contributions of different modules.
+To isolate component effects, all image frames are uploaded, eliminating the influence of frame selection and upload strategies.
+As shown in the Tab.5, the first row corresponds to the baseline where the LVLM and SVLM operate in parallel—using the LVLM output when it is returned in time, and otherwise defaulting to the SVLM.
+This baseline represents the upper-bound performance of existing cloud–edge collaborative VLM architectures.
+Compared to this baseline, we incrementally integrate each proposed module.
+Results show that CRM yields the most significant performance gain.
+Building on CRM, the integration of the other two methods further enhances overall system performance across all three metrics.
+CRM GGVR GRVR Micro-F1 Macro-F1 0-1 EM
+64.36 46.55 22.35
+70.47 55.43 28.85
+70.77 55.58 29.20
+71.15 57.11 29.41
 Table 5: Ablation Study.
 To assess the contribution of each component, we use full-frame upload and LVLM–SVLM parallel inference as the baseline, and evaluate performance improvements as modules are added incrementally.
-SVLM LVLM Micro-F1 Macro-F1 0-1 EM
-Baseline SpotVLM Baseline SpotVLM Baseline SpotVLM
-Qwen2.5-VL-3B Qwen2.5-VL-72B 66.87 73.02 43.07 51.56 24.39 33.17
-MiniCPM-o-2.6-int4 Qwen2.5-VL-72B 67.68 73.56 43.47 49.63 31.38 35.71
-Qwen2.5-VL-3B GPT-4o 66.34 74.25 47.85 53.74 22.53 30.79
-Table 3: Adaptability Study. Testing the SpotVLM architecture with various SVLM–LVLM combinations on nuScenes.
-5.2 Overall Performance
-To evaluate the effectiveness of different architectures, we assess six methods under a Wi-Fi network environment: locally deployed Qwen2.5‑VL‑3B (Bai et al.
-2025), cloud-based Qwen2.5‑VL‑72B (Bai et al.
-2025), Distributed VLM, LAECIPS, ADAS, and SpotVLM.
-The results are presented in Tab.1.
-Due to cloud latency, methods that rely entirely on large models (e.g., Qwen2.5-VL-72B, Distributed VLM) yield lower F1 scores on real-time tasks than collaborative approaches with small models.
-However, the 0–1 EM of Qwen2.5-VL-72B remains higher than that of the small model alone, showing its superior accuracy when responses are timely.
-Additionally, Qwen2.5-VL-3B produces visual features larger than the original image, making Distributed VLM, which transmits features to the cloud, slightly less effective than running full inference on the cloud model.
-Collaborative architectures like LAECIPS and SpotVLM perform consistently well across all four datasets.
-When cloud responses are timely, they effectively leverage large-model outputs, achieving better overall metrics than Qwen2.5-VL-3B.
-ADAS adopts a latency-adaptive offloading strategy.
-In the first three datasets, long outputs from multi-object recognition and dual-image inputs in gesture recognition result in high cloud latency.
-Under these conditions, ADAS relies more on the small model, yielding higher F1 scores but lower 0–1 EM compared to LAECIPS.
-In the fourth dataset, shorter outputs and fewer images per inference lead to more timely cloud responses.
-ADAS thus relies more on the output of the LVLM.
-Meanwhile, LAECIPS depends more on the SVLM for simple samples in this dataset, resulting in slightly lower performance than ADAS.
-To ensure a fair comparison across methods, we recorded the number of cloud service invocations for ADAS and LAECIPS, and used their average values to set the interval for SpotVLM.
-Benefiting from its ability to effectively leverage the high-quality outputs of the LVLM under latency constraints, SpotVLM consistently achieved the best results across all tasks and collaborative strategies.
-To provide a more intuitive comparison, we select one representative example from each of the four datasets for qualitative analysis.
-As shown in the Fig.4, for the real-time multi-object recognition task, all three methods rely on the outputs of the SVLM, with SpotVLM achieving the best performance due to guidance from the historical results of the large model.
-In the gesture recognition task, as two images need to be transmitted and processed simultaneously, the cloud-based large model fails to return valid results in real time, leading to No response for ADAS.
-In contrast, SpotVLM leverages the history-enhanced small model to produce accurate classification results.
-For the video-frame captioning task, both ADAS and SpotVLM adopt the results quickly returned by the large model.
-5.3 Ablation Study
-We conduct a progressive ablation study on the nuScenes dataset to evaluate the individual and combined contributions of Contextual Replacement Module (CRM), Grounding-Guided Visual Representation (GGVR), and Grounding-Tracked Visual Representation (GTVR) in the SpotVLM framework.
-To isolate component effects, all image frames are uploaded, eliminating the influence of frame selection and upload strategies.
-As shown in the Tab.2, the first row corresponds to the baseline where the large and small models operate in parallel—using the LVLM output when it is returned in time, and otherwise defaulting to the SVLM.
-This baseline represents the upper-bound performance of current cloud–edge collaborative VLM architectures.
-Compared to this baseline, we incrementally integrate each proposed module.
-Results show that CRM yields the most significant performance gain.
-GGVR is used to reduce the number of visual tokens, while GTVR guides the SVLM to maintain consistent detection of object categories present in the historical context.
-Building on CRM, the integration of these two methods further enhances overall system performance across all three metrics.
-5.4 Adaptability Study
-To evaluate the effectiveness of the proposed framework under different combinations of model capacities, we conducted additional experiments by replacing both the small and large models.
-As in the ablation study, all image frames are uploaded and processed by the large model, and the upper bound is represented by a parallel inference strategy that always adopts the output of large model when available.
-Specifically, we adopt MiniCPM-o (Yao et al.
-2024) as the SVLM on the edge side and GPT-4o (Hurst et al.
-2024) as the LVLM on the cloud side.
-As shown in the Tab.3, the proposed SpotVLM coordination strategy consistently improves real-time inference performance across model configurations.
-With the continuous advancement of small and large VLMs, more powerful models can be seamlessly integrated into the SpotVLM framework to further enhance system capabilities.
-Figure 5: Latency simulation.
-Manual simulation of total processing latency on the cloud LVLM.
-Red pentagrams denote the performance bounds of LVLM and SVLM; the dashed line shows the SVLM-only performance.
-5.5 Latency Analysis
-To evaluate the impact of cloud latency on SpotVLM, we conduct experiments on the nuScenes dataset using Qwen-2.5-VL-3B as the edge-side SVLM and preserving the outputs of Qwen-2.5-VL-72B as the cloud-side LVLM.
-We manually simulate various latency conditions by setting different return delays for the LVLM responses.
-When the latency is 1 second, which means that the LVLM output always arrives within the real-time threshold, the system achieves performance equivalent to directly using the LVLM (denoted by the star marker).
-Starting from a 2-second delay, the LVLM result can only be incorporated as historical context for guiding the SVLM.
-Even under this condition, SpotVLM significantly outperforms the standalone SVLM (indicated by the dashed line).
-As latency increases, the temporal correlation between frames weakens, leading to a gradual decline in overall performance.
-However, SpotVLM continues to maintain a notable advantage even under a 9-second delay.
-This experiment reveals that the performance upper bound is achieved when the LVLM returns results in real time for every frame, while the lower bound corresponds to using only the SVLM.
-The proposed collaborative architecture offers strong robustness under fluctuating latency conditions.
-Moreover, the results also suggest that cloud-side LVLM invocation can be dynamically scheduled.
-By adjusting the upload interval according to task requirements and performance targets, the system can reduce the frequency of cloud model calls without significantly compromising overall effectiveness.
+4.4.2 Adaptability Study
+To evaluate the effectiveness of the proposed framework under different combinations of model capacities, we conduct additional experiments by replacing both the SVLMs and LVLMs.
+All image frames are uploaded as described in Sec.4.4.1.
+Specifically, we add MiniCPM-o [yao2024minicpm] and VLM-R1 [shen2025vlm] as the SVLMs and GPT-4o [hurst2024gpt] as the LVLM.
+As shown in the Fig.7, the proposed edgeVLM coordination strategy consistently improves real-time inference performance across model configurations.
+VLM-R1 is an SVLM enhanced through distillation and reinforcement learning.
+Experiments demonstrate that powerful models can be seamlessly integrated into the edgeVLM framework to further enhance system capabilities.
 5 Conclusion
 In this work, we propose a novel context transfer paradigm that enables LVLMs to provide real-time guidance to SVLMs through semantic and visual contexts.
-Building on this paradigm, we introduce SpotVLM, a collaborative framework that integrates two corresponding modules: Context Replacement Module and Visual Focus Module.
-Extensive experiments across diverse tasks, datasets, and model combinations demonstrate that SpotVLM outperforms existing collaboration methods.
+Building on this paradigm, we introduce edgeVLM, a collaborative framework that integrates two corresponding modules: Context Replacement and Visual Focus.
+Extensive experiments across diverse tasks, datasets, and model combinations demonstrate that edgeVLM outperforms existing collaboration methods.
 Moving forward, we envision integrating stronger foundation models and adaptive scheduling strategies into the collaborative framework to further optimize performance in real-world deployments.